OpenStudy (dls):

The slope of the tangent to the curve y=x^2 - x at the point where the line y=2 cuts the curve in the first quadrant is?

OpenStudy (dls):

dy/dx=2x-1 2=x^2-x gives x=2 and x=-1 dy/dx=3 and -3 I'm getting 2 options..but the answer is either of them,how do i decide?

OpenStudy (cwrw238):

dy'dx = 3 is in first quadrant

OpenStudy (dls):

why cant it be -3?

OpenStudy (cwrw238):

the point (2,2) is in the first quadrant (the point where y = 2 cuts the curve) the slope at this point is 2(2) - 1 = 3

OpenStudy (dls):

oh,didn't realise this small thing. Thanks!

OpenStudy (cwrw238):

yw